OpenStudy (anonymous):

Identify the discontinuities, graph, and find the zeros of each rational function. f(x) = (4x^2-36x)/(x-9) I don't really understand how to do this at all. Any help is appreciated.

OpenStudy (jdoe0001):

where the function doesn't have a continuous line, that is where it jumps or is undefined like if x = 9, x-9 will make the denominator 0, thus undefined, so, x cannot be = 9

OpenStudy (jdoe0001):

$$ \frac{4x^2-36x}{x-9} \implies \frac{4x(x-9)}{x-9} $$

OpenStudy (jdoe0001):

getting common factor

OpenStudy (jdoe0001):

\(\cfrac{4x(x-9)}{x-9} = 0\)

OpenStudy (jdoe0001):

so, it seems to be the line y = 4x, with a "hole" in x = 9, the zeros, would be the origin

OpenStudy (jdoe0001):